# Effects of dark boson mediated feeble interaction between dark matter (DM) and quark matter on $f$-mode oscillation of DM admixed quark stars

### v1

@@ Abstract
We investigate the behaviour of the prominent non-radial fundamental $f$-mode oscillations of dark matter (DM) admixed strange quark stars (DMSQSs), by adopting an equation of state (EoS) developed in Ref. Sen and Guha (2022), which considers the possible presence of feebly interacting DM in strange quark stars (SQSs) for the first time. Within the model, feeble interaction between fermionic DM $\chi$ and strange quark matter (SQM) is invoked via a vector new physics mediator $\xi$ with coupling strength $y_{\xi}$. The pure SQM is described by the vector MIT Bag model. By varying different EoS parameters, the structural properties viz. the mass, radius and tidal deformability ($\Lambda$) of the DMSQSs are studied with respect to various astrophysical constraints. We study in detail the $f$-mode spectra within the Cowling approximation by obtaining the frequencies as a function of mass, compactness and $\Lambda$ of the star. To the best of our knowledge, this study represents the first analysis of non-radial $f$-mode oscillations of DMSQSs. Our investigation indicates that the presence of DM and its interaction with SQM has great impact on the $f$-modes. We show that the $f$-mode frequencies are larger for DMSQSs, which are largely populated with massive DM fermions, compared to the SQSs. Further, we obtain a linear empirical relation between the $f$-modes and the average density of the star. We also find that the mass-scaled angular frequency varies universally with compactness for DMSQSs.
@@ I Introduction
Many theoretical studies predict that strange quark stars (SQSs), composed of $u$, $d$ and $s$ quarks, may exist Olinto (1987). Such prediction is based on the Bodmer-Witten conjecture, which states that the energy per baryon density $\varepsilon/\rho_{B}$ of strange quark matter (SQM) is lower than that of pure nucleonic system Bodmer (1971); Chin and Kerman (1979); Witten (1984); Farhi and Jaffe (1984). Consequently, over many years, the theoretical modeling of SQM has been a topic of great interest. The first of such kind of models proposed is the well known MIT Bag model Chodos et al. (1974), which is later modified to ensure the repulsive effect between the quarks by introducing a vector meson as mediator (vBag model) Klahn and Fischer (2015); Cierniak et al. (2019); Gomes et al. (2019); Franzon et al. (2016); Wei et al. (2019); Lopes et al. (2021); Kumar et al. (2022). In the MIT Bag model, the difference in energy density between the perturbative vacuum and the true vacuum is represented by a constant $B$ whose value is still not well known. The stability of SQS in terms of $\varepsilon/\rho_{B}$ is found to be controlled by $B$ Farhi and Jaffe (1984); Torres and Menezes (2013) and further, the allowed range of $B$ for different vector meson couplings ($G_{V}$) is estimated in Ref. Lopes et al. (2021).
@@
The compact stars like SQSs are extremely gravitating objects which are capable of accreting matter from their surroundings and such accreted matter may include DM Baryakhtar et al. (2017); Joglekar et al. (2020a); Bell et al. (2019); Joglekar et al. (2020b); Bramante and Raj (2024). Few other phenomena may also be attributed to the possible presence of DM inside compact stars; like, dark decays of neutrons Husain et al. (2022); Husain and Thomas (2023); Husain et al. (2023); Zhou (2023), inheritance of DM from the supernovae Nelson et al. (2019) etc. It is therefore reasonable to consider the DM particle density $\rho_{\chi}$ to be almost constant throughout the radial profile of the star Panotopoulos and Lopes (2017a); Guha and Sen (2021); Sen and Guha (2021, 2022); Guha and Sen (2024). DM thus becomes structural part of the SQSs, consequently allowing us to study the properties of DM admixed SQSs (DMSQSs) in the light of the different observational and astrophysical constraints on the structural properties of compact stars Yang and Pi (2024); Rezaei (2023); Panotopoulos and Lopes (2017b); Zheng and Chen (2016); Mukhopadhyay and Schaffner-Bielich (2016); Jiménez and Fraga (2022); Yang et al. (2021); Mafa Takisa et al. (2020); Panotopoulos and Lopes (2018a, b); Cassing et al. (2023); Ferreira and Fraga (2023). DM may or may not interact with SQM; however, if they interact, then the interaction strength must be extremely weak in order to prevent the star from collapsing Zheng and Chen (2016). Based on this notion, in Ref. Sen and Guha (2022), a feeble interaction between SQM and fermionic DM $\chi$ via a dark vector boson mediator $\xi$ was introduced. The pure SQM is described with the vBag model. The mass of DM fermion $m_{\chi}$ and the mass of vector mediator $m_{\xi}$ are chosen by following the self-interaction constraint $\sigma_{T}/m_{\chi}\leq 1.25~{}\rm{cm^{2}/g}$ Randall et al. (2008); Robertson et al. (2017) from the Bullet cluster Tulin et al. (2013); Tulin and Yu (2018); Hambye and Vanderheyden (2020). Here, $\sigma_{T}$ is the self-scattering transfer cross-section. Further, the coupling $y_{\xi}$ between $\chi$ and $\xi$ are obtained by satisfying the present day relic abundance of DM fermion $\chi$ Belanger et al. (2014); Gondolo and Gelmini (1991); Guha et al. (2019). In the present work, we follow the same treatment as in Ref. Sen and Guha (2022) for the interaction between DM and SQM, in order to investigate its effects on the non-radial oscillations of the DMSQSs.
@@
Gravitational waves (GWs) can be produced by binary neutron star (BNS) mergers as well as from isolated neutron stars (NSs). Following the first detection of a GW signal from a binary black hole merger Abbott et al. (2016), many more signals have been detected from supernovae, NS mergers and supermassive black hole mergers. Furthermore, the NanoGrav collaboration recently revealed the existence of a cosmic gravitational wave background  Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023). The detection of GWs from a BNS merger (GW170817) Abbott et al. (2017) by the advanced LIGO and VIRGO collaborations combined with the multi-messenger observations of BNS systems have paved a new way for studying the interior composition of compact stars. The non-radial pulsations of compact stars are of current interest in the era of gravitational wave detection because they are related to the emission of GWs of different frequencies viz., fundamental ($f$), pressure ($p$), rotational ($r$), space-time ($w$) and gravity ($g$) modes Kokkotas and Schmidt (1999). There have been a lot of studies on these oscillation modes done over the past few years Lindblom and Owen (2002); Jha et al. (2010); Flores and Lugones (2014); Andersson and Pnigouras (2019); Jyothilakshmi et al. (2022); Zhao and Lattimer (2022); Kumar et al. (2024), among which $f$-modes have drawn a great interest since their frequencies ($1-3$ kHz) lie in the range of the future GW detectors such as the Einstein Telescope and the Cosmic Explorer and also the LIGO O4 run. With these next generation GW detectors the $f$-mode oscillations are expected to be observed; which will provide more powerful insights to the composition and interaction of matter inside compact stars and help in constraining the EoS in stringent manner.
@@
For the purpose of gravitational wave asteroseismology, the estimation of the gravitational-radiation flux from compact stars is based on the theory of non-radial oscillations, which was initially developed in the framework of general relativity in Thorne and Campolattaro (1967); Lindblom and Detweiler (1983). Later, with the help of Cowling approximations Cowling (1941), the derivation of the integrated solution was simplified Sotani et al. (2011) by neglecting the metric perturbations. Although, this approximation can result in overestimation of the $f$-modes upto 30% Zhao and Lattimer (2022), it is widely used due to its ability to bring out qualitatively the essential characteristics of $f$-modes. Recently, there have been several studies on non-radial $f$-mode oscillations using the Cowling approximation for different stellar composition Sotani et al. (2011); Flores and Lugones (2014); Ranea-Sandoval et al. (2018); Vásquez Flores et al. (2019); Pradhan et al. (2022); Kalita et al. (2024); Zhang et al. (2024); Jyothilakshmi et al. (2024). Further, the non-radial $f$-mode oscillations of dark matter admixed NSs Das et al. (2021); Hong et al. (2023); Shirke et al. (2024); Sen and Guha (2024) were also studied. In the case of DMSQSs, although the radial oscillations have attracted the interest of the community recently Panotopoulos and Lopes (2017b, 2018a); Jiménez and Fraga (2022); Zhen et al. (2024); the non-radial oscillations have not been explored so far. This sets the motivation for our analysis. In the present work, we are particularly interested to show how the presence of DM and its interaction with SQM affects the non-radial $f$-mode frequency of the oscillation of the DMSQSs using the Cowling approximation.
@@ II.2 Structural properties and non-radial oscillations of dark matter admixed quark stars
The coupled differential equations, Eqs. (13) and (14), are integrated from the center to the surface of the star by assuming an initial value of $\omega^{2}$. After each integration the value of $\omega^{2}$ is improved using Ridder’s method until Eq. (16) is satisfied.
@@ III Results
In this section, we discuss the results obtained for $f$-mode oscillations using the Cowling approximation (given in Sec. II.2) for DMSQSs (given in Sec. II.1). First, we compute the DMSQSs EoS using Eqs. (II.1) and (LABEL:P). With the DMSQSs EoS, we then obtain the static stellar structure by numerically solving the TOV equations given by Eqs. (10) to (12) from the center to the surface of the star. We then numerically integrate the coupled differential Eqs. (13) and (14) from the center to the surface of the star using the appropriate boundary conditions given by Eqs. (15) and (16). The $f$-mode frequencies are obtained with the required precision using Ridder’s method. We also calculate the tidal deformability $\Lambda$ of DM admixed QSs using the Eqs. (18) and (17). In order to study the effect of DM for different values of the SQM vector coupling constant $G_{V}$, we vary the DM particle mass $m_{\chi}$ and the Fermi momentum of the DM fermions ${k_{F}}_{\chi}$ and study their effects on the stellar structure and $f$-mode oscillation of the DMSQSs. We have investigated the explicit dependence of the DMSQS properties on $m_{\chi}$ and ${k_{F}}_{\chi}$ by varying them individually. While varying $m_{\chi}$ and ${k_{F}}_{\chi}$ we have kept the vBag parameters like $G_{V}$ and corresponding value of $B_{av}$ fixed. For example, we fixed $G_{V}$=0.3 fm2 and the corresponding value of $B_{av}$ when we varied $m_{\chi}$. Again we fix $G_{V}$=0.5 fm2 with the corresponding value of $B_{av}$ while varying ${k_{F}}_{\chi}$. While varying $m_{\chi}$, we have fixed ${k_{F}}_{\chi}$ and vice versa. This way, the dependence of the properties of DMSQS on both the mass and the number density of DM are examined explicitly and individually in the present paper. While varying the parameters of dark sector, the parameters of the MIT vbag model are kept fixed in order to study the distinct role that DM plays on the properties of DMSQS.
@@
Finally, for different parameterizations of DMSQSs, we obtain a linear (empirical) relation between the $f$-mode frequency and the average density $\bar{\rho}=\sqrt{\bar{M}/\bar{R}^{3}}$, where $\bar{M}=M/(1.4$ $M_{\odot}$) and $\bar{R}=R/(10$ km). The empirical formula of the form, $f=a+b\bar{\rho}$ was first obtained by Andersson and Kokkotas Andersson and Kokkotas (1996) and by knowing the values of $f$-mode frequencies, we can calculate the mass and radius of the star. This could therefore help in constraining the EoS more closely Andersson and Kokkotas (1996). For our EoSs, we obtain the values of $a$ and $b$ as 1.367 kHz and 0.87 kHz respectively. In Fig. 9(a), we plot the relation obtained for DMSQSs along with the empirical relations from relevant earlier works for stellar models with dark matter admixed hyperon stars (DKBP) Das et al. (2021) and hybrid stars (BFG) Benhar et al. (2004). We find that our fit relation differ considerably from the already existing relations from DKBP and BFG. However, our fit is found to be more close to the fit for hybrid stars denoted by BFG. Further, we also obtain the correlation between the angular frequency scaled by mass ($\omega M$) and the compactness ($C$) of the DMSQSs. In Ref. Wen et al. (2019), Wen et al. found that the variation of $\omega M$ as a function of $C$ is universal for $f$-modes. Similarly, in Fig. 9(b), we fit the mass scaled $\omega$ with $C$ for different parameterizations of DMSQSs. The universal relation was found to take the form $\omega M$ (kHz km) $=129.50\,(M/R)-3.48$, where the coefficients are obtained after performing a linear fit. In order to obtain these correlations we varied the values of $G_{V}$ from 0.3 to 0.7 GeV and $m_{\chi}$ from 50 to 100 GeV.

### v2

@@ Abstract
We investigate the behavior of the prominent non-radial fundamental $f$-mode oscillations of dark matter (DM) admixed strange quark stars (DMSQSs), by adopting an equation of state (EoS) developed in Ref. Sen and Guha (2022), which considers the possible presence of feebly interacting DM in strange quark stars (SQSs) for the first time. Within the model, feeble interaction between fermionic DM $\chi$ and strange quark matter (SQM) is invoked via a vector new physics mediator $\xi$ with coupling strength $y_{\xi}$. The pure SQM is described by the vector MIT Bag model. By varying different EoS parameters, the structural properties viz. the mass, radius and tidal deformability ($\Lambda$) of the DMSQSs are studied with respect to various astrophysical constraints. We study in detail the $f$-mode spectra within the Cowling approximation by obtaining the frequencies as a function of mass, compactness and $\Lambda$ of the star. To the best of our knowledge, this study represents the first analysis of non-radial $f$-mode oscillations of DMSQSs. Our investigation indicates that the presence of DM and its interaction with SQM has great impact on the $f$-modes. We show that the $f$-mode frequencies are larger for DMSQSs, which are largely populated with massive DM fermions, compared to the SQSs. Further, we obtain a linear empirical relation between the $f$-modes and the average density of the star. We also find that the mass-scaled angular frequency varies universally with compactness and $\ln{\Lambda}$ for DMSQSs. Further, our studies indicate that the inclusion of DM in compact stars reduces the deviation of $f$-mode frequency from general relativistic to Cowling approximation.
@@ I Introduction
Many theoretical studies predict that strange quark stars (SQSs), composed of $u$, $d$ and $s$ quarks, may exist Olinto (1987); Farhi and Jaffe (1984); Torres and Menezes (2013); Miao et al. (2021); Li et al. (2021); Lugones and Grunfeld (2023); Li et al. (2020, 2019); Luo et al. (2019); Pal and Chaudhuri (2023). Such prediction is based on the Bodmer-Witten conjecture, which states that the energy per baryon density $\varepsilon/\rho_{B}$ of strange quark matter (SQM) is lower than that of pure nucleonic system Bodmer (1971); Chin and Kerman (1979); Witten (1984); Farhi and Jaffe (1984). Consequently, over many years, the theoretical modeling of SQM has been a topic of great interest. The first of such kind of models proposed is the well known MIT Bag model Chodos et al. (1974), which is later modified to ensure the repulsive effect between the quarks by introducing a vector meson as mediator (vBag model) Klahn and Fischer (2015); Cierniak et al. (2019); Gomes et al. (2019); Franzon et al. (2016); Wei et al. (2019); Lopes et al. (2021); Kumar et al. (2022). In the MIT Bag model, the difference in energy density between the perturbative vacuum and the true vacuum is represented by a constant $B$ whose value is still not well known. The stability of SQS in terms of $\varepsilon/\rho_{B}$ is found to be controlled by $B$ Farhi and Jaffe (1984); Torres and Menezes (2013) and further, the allowed range of $B$ for different vector meson couplings ($G_{V}$) is estimated in Ref. Lopes et al. (2021).
@@
The compact stars like SQSs are extremely gravitating objects which are capable of accreting matter from their surroundings and such accreted matter may include DM Baryakhtar et al. (2017); Joglekar et al. (2020a); Bell et al. (2019); Joglekar et al. (2020b); Bramante and Raj (2024). Few other phenomena may also be attributed to the possible presence of DM inside compact stars; like, dark decays of neutrons Husain et al. (2022); Husain and Thomas (2023); Husain et al. (2023); Zhou (2023), inheritance of DM from the supernovae Nelson et al. (2019) etc. It is therefore reasonable to consider the DM particle density $\rho_{\chi}$ to be almost constant throughout the radial profile of the star Panotopoulos and Lopes (2017a); Guha and Sen (2021); Sen and Guha (2021, 2022); Guha and Sen (2024). The compact objects with DM admixture may also be studied within the two fluid approach and several works are there in this direction Leung et al. (2012); Rezaei (2017); Mukhopadhyay et al. (2017); Das et al. (2022); Mariani et al. (2024); Barbat et al. (2024); Zhen et al. (2024); Thakur et al. (2024), though the interaction between DM and nuclear matter is not considered.
@@
DM thus becomes structural part of the SQSs, consequently allowing us to study the properties of DM admixed SQSs (DMSQSs) in the light of the different observational and astrophysical constraints on the structural properties of compact stars Yang and Pi (2024); Rezaei (2023); Panotopoulos and Lopes (2017b); Zheng and Chen (2016); Mukhopadhyay and Schaffner-Bielich (2016); Jiménez and Fraga (2022); Yang et al. (2021); Mafa Takisa et al. (2020); Panotopoulos and Lopes (2018a, b); Cassing et al. (2023); Ferreira and Fraga (2023). DM may or may not interact with SQM; however, if they interact, then the interaction strength must be extremely weak in order to prevent the star from collapsing Zheng and Chen (2016). Based on this notion, in Ref. Sen and Guha (2022), a feeble interaction between SQM and fermionic DM $\chi$ via a dark vector boson mediator $\xi$ was introduced. The pure SQM is described with the vBag model. The mass of DM fermion $m_{\chi}$ and the mass of vector mediator $m_{\xi}$ are chosen by following the self-interaction constraint $\sigma_{T}/m_{\chi}\leq 1.25~{}\rm{cm^{2}/g}$ Randall et al. (2008); Robertson et al. (2017) from the Bullet cluster Tulin et al. (2013); Tulin and Yu (2018); Hambye and Vanderheyden (2020). Here, $\sigma_{T}$ is the self-scattering transfer cross-section. Further, the coupling $y_{\xi}$ between $\chi$ and $\xi$ are obtained by satisfying the present day relic abundance of DM fermion $\chi$ Belanger et al. (2014); Gondolo and Gelmini (1991); Guha et al. (2019). In the present work, we follow the same treatment as in Ref. Sen and Guha (2022) for the interaction between DM and SQM, in order to investigate its effects on the non-radial oscillations of the DMSQSs.
@@
Gravitational waves (GWs) can be produced by binary neutron star (BNS) mergers as well as from isolated neutron stars (NSs). Following the first detection of a GW signal from a binary black hole merger Abbott et al. (2016), many more signals have been detected from supernovae, NS mergers and supermassive black hole mergers. Furthermore, the NanoGrav collaboration recently revealed the existence of a cosmic gravitational wave background  Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023). The detection of GWs from a BNS merger (GW170817) Abbott et al. (2017) by the advanced LIGO and VIRGO collaborations combined with the multi-messenger observations of BNS systems have paved a new way for studying the interior composition of compact stars. The non-radial pulsations of compact stars are of current interest in the era of gravitational wave detection because they are related to the emission of GWs of different frequencies viz., fundamental ($f$), pressure ($p$), rotational ($r$), space-time ($w$) and gravity ($g$) modes Kokkotas and Schmidt (1999). There have been a lot of studies on these oscillation modes done over the past few years Lindblom and Owen (2002); Jha et al. (2010); Flores and Lugones (2014); Andersson and Pnigouras (2019); Jyothilakshmi et al. (2022); Zhao and Lattimer (2022); Kumar et al. (2024a); Guha et al. (2024), among which $f$-modes have drawn a great interest since their frequencies ($1-3$ kHz) lie in the range of the future GW detectors such as the Einstein Telescope and the Cosmic Explorer and also the LIGO O4 run. With these next generation GW detectors the $f$-mode oscillations are expected to be observed; which will provide more powerful insights to the composition and interaction of matter inside compact stars and help in constraining the EoS in stringent manner.
@@
For the purpose of gravitational wave asteroseismology, the estimation of the gravitational-radiation flux from compact stars is based on the theory of non-radial oscillations, which was initially developed in the framework of general relativity in Thorne and Campolattaro (1967); Lindblom and Detweiler (1983). Later, with the help of Cowling approximations Cowling (1941), the derivation of the integrated solution was simplified Sotani et al. (2011) by neglecting the metric perturbations. Although, this approximation can result in overestimation of the $f$-modes upto 30% Zhao and Lattimer (2022) compared to a fully general relativistic treatment preferred for a precise quantitative analysis; Cowling approximation is widely used to study the characteristics of $f$-modes. In the present work, we calculated the deviation of the $f$-mode frequencies obtained using the Cowling approximation from the fully relativistic approach for neutron star (NS), dark matter admixed neutron star (DMNS) and SQS. We found that the deviation percentage of DMNS is less compared to that of NS and SQS, indicating that inclusion of DM may lower the deviation. Thus, we believe that using the Cowling approximation in our analysis of DMSQSs may not result in a higher deviation percentage.
Recently, there have been several studies on non-radial $f$-mode oscillations using the Cowling approximation for different stellar composition Sotani et al. (2011); Flores and Lugones (2014); Ranea-Sandoval et al. (2018); Vásquez Flores et al. (2019); Pradhan et al. (2022); Kalita et al. (2024); Zhang et al. (2024); Jyothilakshmi et al. (2024, 2025a, 2025b). Further, the non-radial $f$-mode oscillations of quark stars Cutler et al. (1990); Kojima and Sakata (2002); Sotani and Harada (2003) and dark matter admixed NSs Das et al. (2021); Hong et al. (2023); Shirke et al. (2024); Sen and Guha (2024) were also studied. In the case of DMSQSs, although the radial oscillations have attracted the interest of the community recently Panotopoulos and Lopes (2017b, 2018a); Jiménez and Fraga (2022); Zhen et al. (2024); the non-radial oscillations have not been explored so far. This sets the motivation for our analysis.
In the present work, we are particularly interested to show how the presence of DM and its interaction with SQM affects the non-radial $f$-mode frequency of the oscillation of the DMSQSs using the Cowling approximation.
@@ II.2 Structural properties and non-radial oscillations of dark matter admixed quark stars
The coupled differential equations, Eqs. (13) and (14), are integrated from the center to the surface of the star by assuming an initial value of $\omega^{2}$. After each integration the value of $\omega^{2}$ is improved using Ridders’ method until Eq. (16) is satisfied.
@@ III Results
In this section, we discuss the results obtained for $f$-mode oscillations using the Cowling approximation (given in Sec. II.2) for DMSQSs (given in Sec. II.1). First, we compute the DMSQSs EoS using Eqs. (II.1) and (LABEL:P). With the DMSQSs EoS, we then obtain the static stellar structure by numerically solving the TOV equations given by Eqs. (10) to (12) from the center to the surface of the star. We then numerically integrate the coupled differential Eqs. (13) and (14) from the center to the surface of the star using the appropriate boundary conditions given by Eqs. (15) and (16). The $f$-mode frequencies are obtained with the required precision using Ridders’ method. We also calculate the tidal deformability $\Lambda$ of DM admixed QSs using the Eqs. (18) and (17). In order to study the effect of DM for different values of the SQM vector coupling constant $G_{V}$, we vary the DM particle mass $m_{\chi}$ and the Fermi momentum of the DM fermions ${k_{F}}_{\chi}$ and study their effects on the stellar structure and $f$-mode oscillation of the DMSQSs. We have investigated the explicit dependence of the DMSQS properties on $m_{\chi}$ and ${k_{F}}_{\chi}$ by varying them individually. While varying $m_{\chi}$ and ${k_{F}}_{\chi}$ we have kept the vBag parameters like $G_{V}$ and corresponding value of $B_{av}$ fixed. For example, we fixed $G_{V}$=0.3 fm2 and the corresponding value of $B_{av}$ when we varied $m_{\chi}$. Again we fix $G_{V}$=0.5 fm2 with the corresponding value of $B_{av}$ while varying ${k_{F}}_{\chi}$. While varying $m_{\chi}$, we have fixed ${k_{F}}_{\chi}$ and vice versa. This way, the dependence of the properties of DMSQS on both the mass and the number density of DM are examined explicitly and individually in the present paper. While varying the parameters of dark sector, the parameters of the MIT vbag model are kept fixed in order to study the distinct role that DM plays on the properties of DMSQS.
@@
Finally, for different parameterizations of DMSQSs, we obtain a linear (empirical) relation between the $f$-mode frequency and the average density $\bar{\rho}=\sqrt{\bar{M}/\bar{R}^{3}}$, where $\bar{M}=M/(1.4$ $M_{\odot}$) and $\bar{R}=R/(10$ km). The empirical formula of the form, $f=a+b\bar{\rho}$ was first obtained by Andersson and Kokkotas Andersson and Kokkotas (1996) and by knowing the values of $f$-mode frequencies, we can calculate the mass and radius of the star. This could therefore help in constraining the EoS more closely Andersson and Kokkotas (1996). For our EoSs, we obtain the values of $a$ and $b$ as 1.367 kHz and 0.87 kHz respectively. In Fig. 9(a), we plot the relation obtained for DMSQSs along with the empirical relations from relevant earlier works for stellar models with dark matter admixed hyperon stars (DKBP) Das et al. (2021) and hybrid stars (BFG) Benhar et al. (2004). We find that our fit relation differ considerably from the already existing relations from DKBP and BFG. However, our fit is found to be more close to the fit for hybrid stars denoted by BFG. Additionally, there are other model independent relations that provide a stronger correlation between the $f$-mode frequency, and both compactness and tidal deformability. First, we obtain the correlation between the angular frequency scaled by mass ($\omega M$) and the compactness ($C$) of the DMSQSs. In Ref. Wen et al. (2019), Wen et al. found that the variation of $\omega M$ as a function of $C$ is universal for $f$-modes. Similarly, in Fig. 9(b), we fit the mass scaled $\omega$ with $C$ for different parameterizations of DMSQSs. The universal relation was found to take the form $\omega M$ $=0.634\,(M/R)-0.0176$, where the coefficients are obtained after performing a linear fit. In order to obtain these correlations we varied the values of $G_{V}$ from 0.3 to 0.7 GeV and $m_{\chi}$ from 50 to 100 GeV.
Figure 10: Mass scaled angular velocity ($\omega M$) as a function of tidal deformability ($\ln{\Lambda}$) of DMSQSs. The solid red line represents the correlation obtained by including all parameterizations of DMSQSs used in the paper. A correlation between $f$-mode frequency and $\Lambda$ obtained by Kumar et al. Kumar et al. (2024b) is also shown.
Further, one can also obtain a relation between the mass-scaled angular velocity and tidal deformability, where $\omega M$ is found to vary universally with $\Lambda$ and the relation is given by Chan et al. (2014); Sotani and Kumar (2021); Pradhan et al. (2022); Shirke et al. (2024)
$\omega M=\sum\limits_{i}\alpha_{i}(\ln{\Lambda})^{i}.$ (19)
In Fig. 10, we plot $\omega M$ as a function of $\Lambda$ for different parameterization of DMSQSs. The correlation obtained is $0.191+0.004(\ln{\Lambda})-6.15\times 10^{-3}(\ln{\Lambda})^{2}+6.56\times 10^{-4}(\ln{\Lambda})^{3}-2.694\times 10^{-5}(\ln{\Lambda})^{4}+3.905\times 10^{-7}(\ln{\Lambda})^{5}$. We have also plotted the relation obtained between $\omega M$ and $\ln{\Lambda}$ given in Ref. Kumar et al. (2024b), using the Cowling approximation: $0.198-2.612\times 10^{-3}(\ln{\Lambda})-4.899\times 10^{-3}(\ln{\Lambda})^{2}+5.547\times 10^{-4}(\ln{\Lambda})^{3}-2.246\times 10^{-5}(\ln{\Lambda})^{4}+2.754\times 10^{-7}(\ln{\Lambda})^{5}$.
We find that both results are in good agreement. Some of the previous works on stars containing quark matter have also obtained the relation between $\omega M$ and $\Lambda$: $\omega M=0.182-6.11\times 10^{-3}(\ln{\Lambda})-4.594\times 10^{-3}(\ln{\Lambda})^{2}+6.066\times 10^{-4}(\ln{\Lambda})^{3}-2.614\times 10^{-5}(\ln{\Lambda})^{4}+2.228\times 10^{-7}(\ln{\Lambda})^{5}$ Pradhan et al. (2024) and $\log(2\pi f/\Lambda)=1.052-0.984\,\log(M_{1.4}/\Lambda)-0.030\,(\log(M_{1.4}/\Lambda))^{2}$ Ranea-Sandoval et al. (2022).
Figure 11: Percentage deviation of $f$-mode frequencies obtained using the Cowling approximation from the fully relativistic treatment for SQS with coupling constant $G_{V}=0.3$ fm2 Lopes et al. (2024) and dark matter admixed neutron star (DMNS) with self-interaction strength $G=100$ fm2 Shirke et al. (2024).
[…]